The average number of pages in a novel is 326 with a standard deviation of 24 pages. If a sample of 50 novels is randomly chosen, what is the probability that the average number of pages in these books is between 319 and 331? 


1
Expert's answer
2021-07-19T14:36:38-0400

"\mu= 326 \\
\sigma= 24 \\
n=50 \\
P(319< \bar{x}<331) = P(\bar{x}<331) -P(\bar{x}<319) \\
=P(Z< \frac{331-326}{24/ \sqrt{50}}) -P(Z< \frac{319-326}{24/ \sqrt{50}}) \\
= P(Z<1.473) -P(Z< -2.062) \\
= 0.9296-0.0196 \\
= 0.91"